### v1

# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

Abstract
This paper introduces misinfo-general, a benchmark dataset for evaluating misinformation models’ ability to perform out-of-distribution generalisation. Misinformation changes rapidly, much quicker than moderators can annotate at scale, resulting in a shift between the training and inference data distributions. As a result, misinformation models need to be able to perform out-of-distribution generalisation, an understudied problem in existing datasets. We identify 6 axes of generalisation—time, event, topic, publisher, political bias, misinformation type—and design evaluation procedures for each. We also analyse some baseline models, highlighting how these fail important desiderata. We make this dataset and accompanying code publicly available [^1] https://github.com/ioverho/misinfo-general
[…]
booktabs
Yesterday’s News: Benchmarking Multi-Dimensional
Out-of-Distribution Generalisation of Misinformation Detection Models
Ivo Verhoeven, Pushkar Mishra, Ekaterina Shutova $\dagger$ ILLC, University of Amsterdam, The Netherlands $\ddagger$ Meta AI, London, United Kingdom i.o.verhoeven@uva.nl, pm576@cantab.ac.uk, e.shutova@uva.nl
1 Introduction
The field of misinformation detection aims to develop classification models that can reliably moderate online content. Despite burgeoning academic interest Wu et al. (2019); Zhou and Zafarani (2020) and impressive classification results on existing datasets, content with the explicit intent to deceive continues to propagate and cause harm.
The rapid evolution of online content, which significantly outpaces model development cycles, partially explains this. For example, recent emerging news events Nair (2024); Reuters (2024), containing novel entities or entity relations, spawned a deluge of misinformation resulting in detrimental societal effects. Given that domain expert labels are prohibitively expensive to procure, in such cases training datasets will lag behind the news landscape. As a result, the inference-time text distribution will quickly shift away from the training data. For misinformation detection to be successful during deployment, models need to be robust to distribution shifts—to mitigate future harms, models must leverage general rules learned from yesterday’s news.
Currently, this property of Out-of-Distribution (OoD) generalisation is lacking in many SoTA models. Performance degrades when evaluated on unseen time periods Bozarth and Budak (2020); Horne et al. (2020b); Kochkina et al. (2023); Stepanova and Ross (2023), events Lee et al. (2021); Ding et al. (2022), domains Hoy and Koulouri (2022); Kochkina et al. (2023); Verhoeven et al. (2024), cultures, or languages Horne et al. (2020a); Chu et al. (2021); Ozcelik et al. (2023).
We partially attribute this to the state of existing datasets. While plentiful, these are usually small, have been collected over a short time-span, are centred around specific events, or contain a homogeneous set of publishers. To learn generalisation, however, modern NLP models require large, diverse pre-training datasets, especially when text or labels are noisy.
This paper presents a novel benchmark dataset that enables training and evaluation for OoD generalisation in misinformation detectors. It balances article quantity and publisher diversity with labelling quality by leveraging rich publisher-level metadata annotations. We identify six axes along which we test generalisation performance, representing the most comprehensive evaluation at scale: 1) time, 2) specific events, 3) topics, 4) publisher style, 5) political bias and 6) misinformation type.
We train a baseline model and provide in-depth analysis of its performance to showcase this dataset’s features. We show generalisation to different classes of publishers to be particularly challenging. Further analysis using the available publisher-level metadata highlights some undesirable model behaviours, such as worse performance on less frequent publishers, and worse classification on different political biases. Finally, we find some initial evidence that the scale and diversity of this dataset can benefit generalisation ability.
2 Related Work
2.1 Generalisable Misinformation Detection
Generalisation abilities of misinformation classifiers have been tested at smaller scales. Horne et al. (2020b) found that performance degrades quickly when evaluating on future events, which Bozarth and Budak (2020) corroborate and extend to changes in domain. The same issues have also been reported in misinformation detection in other modalities Stepanova and Ross (2023); Verhoeven et al. (2024).
Results on existing benchmark datasets are generally not indicative of down-stream performance. Kochkina et al. (2023) found that performance within one dataset vastly overestimates performance on other datasets or time spans. Even when controlling for the time period or topic, Hoy and Koulouri (2022) found that models overfit to the training dataset and perform worse on similar but unseen datasets. In a recent systematic review of the literature, Xiao and Mayer (2024) come to the conclusion that “…detection tasks are often meaningfully distinct from the challenges that online services actually face. Datasets and model evaluation are often non-representative of real-world contexts, and evaluation frequently is not independent of model training” (p. 1).
There has been some interest in developing generalisable or adaptive misinformation detection techniques. This has been attempted through multitask training Lee et al. (2021), utilising external agents Ding et al. (2022); Mosallanezhad et al. (2022), data resampling or active learning Hu et al. (2023), adversarial learning Lin et al. (2022), or gradient-based meta-learning Yue et al. (2022). While these research directions are promising, their utility for out-of-distribution misinformation detection has not been sufficiently tested on large, diverse benchmark data.
2.2 Misinformation Datasets
There exist more misinformation datasets than we can exhaustively list here. Instead, we highlight the $4$ predominant annotation schemes, from most fine-grained to most coarse-grained.
  1. 1.
    Span: experts manually fact-check individual (but complete) statements Wang (2017); Martino et al. (2020); Poldvere et al. (2023)
  2. 2.
    Document: experts manually check the overall veracity of entire documents Potthast et al. (2017); Shu et al. (2019)
  3. 3.
    Claim: experts verify claims related to specific events, while article labels are generated by proxy from their stance to these claims Nielsen and McConville (2022)
  4. 4.
    Publisher: experts label publishers by their propensity for factual reporting. Article labels are generated by proxy through the label of their publisher Horne et al. (2018); Barrón-Cedeño et al. (2019); Cui and Lee (2020)
Fine-grained annotation methods yield high-quality labels, but are prohibitively expensive to procure. As a result, such datasets usually sample from few publishers, or a relatively limited time-span. Using publisher labels as a proxy can introduce noise, as it assumes all publishers’ articles share a label, but is far cheaper. This can enable large collections of articles, with diverse sets of publishers and events. The ‘Claim’ annotation level, is a relatively new method, and could balance the precision of ‘Span’ annotations with the affordability of ‘Publisher’ labels. misinfo-general is an example of ‘Publisher’ labelling.
We recommend Xiao and Mayer (2024) for a more comprehensive review of datasets at the article and claims labelling scales.
2.2.1 Metadata Provenance
For publisher-level annotations and metadata, we scraped Media Bias/Fact Check [^2] https://mediabiasfactcheck.com/ (MBFC). MBFC is a curated database of news publishers, with thorough analyses of publisher origins, bias, and credibility. Despite being run by lay volunteers, MBFC labels correlate well with professional fact-checking sources Kiesel et al. (2019); Broniatowski et al. (2022); Pratelli and Petrocchi (2022). MBFC labels have been used in many earlier works Rashkin et al. (2017); Baly et al. (2018a, 2020a); Burdisso et al. (2024); Casavantes et al. (2024); Szwoch et al. (2024). MBFC labels are dynamic and annotations can change (although this is infrequent). We use the latest available version of labels (Oct. 2024).
2.3 Publisher Reliability Estimation
The estimation of a publisher’s reliability is a relatively well-studied problem. This is usually achieved through a mix of content-based Rashkin et al. (2017); Bianchi et al. (2024) and metadata features Baly et al. (2018a, b, 2019, 2020a, 2020b); Nakov et al. (2021).
Relative to purely content-based classifiers, publisher-level classification can greatly reduce the computational cost needed for classification Burdisso et al. (2024). However, this usually involves incorporating additional historical context, world-knowledge Yang and Menczer (2024) or social context Pratelli et al. (2024).
This mimics how moderators or users might analyse the reliability of a publisher, potentially before ingesting the article contents. However, such approaches might fail in cases of where publishers are unknown, ambiguous or have changed their editorial bias. In those situations, moderation decisions at the article-level is necessary. While misinfo-general is suited for either approach, we focus on testing the generalisation of inductive article-level classifiers.
3 Data
All raw articles come from the various News Landscape (NELA) corpora produced by the MELA lab [^3] https://melalab.github.io (2018; 2019; 2020; 2021; 2023). The corpora cover 2017–2022 (6 iterations) almost continuously, with articles from a diverse group of publishers. In their original form, the 6 iterations together consist of $7.2$ million long-form articles.
The original authors’ goal was to study the dynamic behaviour of news and news publishers. Existing corpora were deemed inadequate due to having 1) few, relatively homogeneous articles or publishers, 2) too narrow a focus on specific events, 3) bias towards popular publishers, and 4) limited ground truth labelling Horne et al. (2018); Norregaard et al. (2019).
3.1 Publisher Labelling
From the 2018 iteration onwards, the NELA datasets come with publisher-level labels. However, due to inconsistencies across dataset iterations and the frequency of labelling errors, we chose to relabel the dataset completely. We first mapped all articles to a consistent set of publishers using the URL domain of the scraped articles before removing any publishers known to be news aggregators or social media sites. After updating the article-publisher mapping, we label each publisher using a more recent version of the MBFC database. We provide further detail in Appendix B.1. Ultimately, we identified $488$ distinct publishers, many of which were falsely attributed in NELA’s original set of publishers. The metadata available for each publisher is provided in Appendix B.6.
3.2 Data Processing
Beyond errors in the article-publisher and publisher-label mappings, the texts themselves frequently contain duplicates or scraping errors. Of the $6.7$M re-labelled articles, roughly $\approx 22\%$ or $1.5$M articles were duplicates. Many of the remaining unique articles were deemed malformed or semantically void. These contain either very little text, substantial amounts of markup or include too many special tokens to be human-readable. We filter these using a few simple rules (see Appendices B.2 and B.3). Altogether, we remove approximately $\approx 43\%$ of all downloaded articles. The final dataset contains $4.2$ million cleaned articles.
In the remaining texts, we mask various forms of private or identifiable information (PII), both to enhance safety and reduce the number of available classification ‘shortcuts’. We furthermore standardize the copyright masking procedure introduced in Gruppi et al. (2020). This introduces $4$ new special tokens: <copyright> replacing NELA’s repeated @ tokens, <twitter>, <url> and <selfref> for any self-references.
Despite our efforts, the datasets retain a level of ‘noise’ customary to data sourced from the internet. For example, articles from the same publisher tend to contain unique by-lines, attribution messages, or donation requests. Further cleaning efforts might reduce the realism of the benchmark.
3.3 Topic Clustering
One of our aims is to test model generalisation across different events and topics. To discover these, we used a modified variant of BERTopic Grootendorst (2022) with a gte-large [^4] https://huggingface.co/thenlper/gte-large Li et al. (2023) backbone. This produced thousands of event clusters for every dataset iteration, each with a TF-IDF representation vector. We aggregate these events into overarching topics by applying spectral clustering to the adjacency matrix induced by the inter-event cosine similarity of the TF-IDF matrix. We arbitrarily limit the number of topics to 10, each with varying numbers of events in them. This process is further described in Appendix B.4.
This extends the work of Litterer et al. (2023) on identifying ‘news storms’ in the NELA corpora to a larger time-span, and a larger set of publishers.
4 Generalisation Taxonomy
Generalisation Axis In Distribution Out-of-Distribution
Time Time
CNN AP Vox
2018 2018 2018
CNN AP Vox
2017 2020 2019
Content Event not COVID-19 events COVID-19 events
[…]
Publisher Publisher
CNN MSNBC OANN
Reuters AP True Activist
[…]
AP Reuters Fox News
Centre Centre Right
Vox Daily Beast True Activist
Left Left Left
[…]
Vox NYT OANN
Reliable Reliable Questionable
MSNBC 911Truth.org Age of Autism
Reliable Conspiracy Pseudosci.
Table 1: A schematic overview of the generalisation taxonomy. The left columns provide relevant generalisation category and axis, whereas the right columns provide examples of in domain and out-of-distribution article sets.
Prior work has found misinformation models to be brittle to various forms of distribution shift. We partially attribute this to a lack of diversity in article content and publishers in the training data. By relaxing the assumption of article-level labelling, we can construct a dataset that is both large and diverse, enabling generalisation-focused training and evaluation across various axes. Each such axis tests performance with respect to a specific shift in the training and inference time input distributions of misinformation detectors.
This paper considers 6 specific generalisation axes. Time-based generalisation measures the extent to which changes in publisher style affect a model’s predictions. Evolution of article content will also impact performance. We focus on two specific forms of such change: 1) due to spontaneous events, which we define as news-worthy happenings with a definite time-span, or 2) due to changing topics, which we define as large, overarching collections of events that remain relatively static over a period measured in years or decades. We expect language to be noticeably different across different collections of events or topics.
The distribution of publishers is also expected to change. All publishers exhibit some editorial bias or style which can be memorized by classification models. One, usually implicit, expectation of misinformation detectors is generalisation to different political biases or misinformation types. Predictions ought to be based on a publisher’s intent, not their norms and values.
This taxonomy is by no means exhaustive, and we discuss some additional categories of domain shift in Sec. 8.
4.1 Data Splits
To operationalise these generalisation axes, we build 6 (+1 baseline) train/test splits of the dataset using the publisher-level metadata available to us. In each scenario, a validation split is sampled i.i.d. from the training set. Throughout, we approximate the same $70$/$10$/$20$% article proportions per training/validation/test split, respectively. We repeat each experiment for each year, independently, 6 times in total. The only exception is the ‘Event’ axis, for which we combine all years into a single dataset.
These splits are carefully constructed to test each form of generalisation in isolation. From the model’s perspective, each experiment seems identical, with similar article counts and class proportions. Briefly, we use the following splits:
  1. 0.
    Uniform: standard stratified random splitting of articles into disjoint sets
Time: the training set comprises a single dataset year, while the evaluation set consists of articles from seen publishers in all other years
Event: we reserve all articles which include keywords related to the COVID-19 pandemic for testing, and train on the rest
Topic: we reserve the smallest topic clusters for the test set, and train on articles from the rest
Publisher: we reserve the least frequent publishers for the test set
Political Bias: we reserve all articles from either all leftist or rightist publishers for testing
Misinformation Type: we reserve all articles from either all ‘Questionable Source’ or ‘Conspiracy-Pseudoscience’ publishers for testing
We include an expanded description of each split’s construction in Appendix A.
Generalisation Form MCC F1 Reliable F1 Unreliable
ID OoD $\Delta$ ID OoD $\Delta$ ID OoD $\Delta$
Uniform 0.46 0.46 0.00 0.86 0.86 0.00 0.57 0.57 0.00
Time 0.46 0.33 -0.13 N/A
Event 0.43 0.46 0.03 0.87 0.86 -0.01 0.52 0.55 0.03
Topic 0.46 0.38 -0.08 0.87 0.84 -0.03 0.56 0.50 -0.06
Publisher 0.48 0.37 -0.10 0.87 0.84 -0.03 0.58 0.53 -0.05
Political Bias
Left 0.49 0.30 -0.19 0.85 0.87 0.02 0.61 0.38 -0.23
Right 0.56 0.19 -0.37 0.95 0.60 -0.34 0.58 0.26 -0.32
Misinformation Type
Conspiracy-Pseudoscience 0.43 0.42 -0.01 0.87 0.82 -0.05 0.53 0.53 0.01
Questionable Source 0.43 0.23 -0.20 0.94 0.62 -0.33 0.41 0.25 -0.16
Table 2: Article-level classification performance for various generalisation axes. Specifically, compares performance on the uniformly split in-domain data (ID) to that of the out-of-distribution evaluation set (OoD). The top row uses uniform splitting for both (OoD is actually ID), serving as a baseline value. We have opted to exclude F1 values for the ’Time’ based splitting, as these compare across datasets with strongly varying class proportions.
5 Experiments and Results
To showcase the utility of misinfo-general for model training and evaluation, we trained a simple yet powerful baseline model. Specifically, we fine-tune an instance of DeBERTa-v3555https://huggingface.co/microsoft/deberta-v3-base He et al. (2023) where we reset the pooler and classification weights but freeze the model’s remaining weights. To enable using dataset-specific tokens, we allow the token embedding layer to train with a very low learning rate. The model’s pre-training data included a closed-source news dataset (CC-News), dated between September 2016 and February 2019 Liu et al. (2019).
We fine-tune the models on different splits outlined in Sec. 4. The validation set is always sampled uniformly at random from the training data. We keep the hyperparameters and compute budget constant (which were tuned on the validation sets of the ‘Uniform’ splits). We outline hyperparameters in Appendix D. Training occurs at the article level. We binarise the article publisher’s MBFC label for training labels, which we optimise for using the cross-entropy loss. All ‘Questionable Source’, ‘Conspiracy-Pseudoscience’, and ‘Satire’ publishers were deemed unreliable-, and all others reliable+. We leave testing other publisher-label mappings to future work.
To assess model performance at the article level, we employ the F1-score computed independently for each class along with the Matthews Correlation Coefficient (MCC). The F1-score’s interpretation is largely dependent on the class proportion, making it less suited to comparison across experiments. MCC is more robust to this.
5.1 OoD Generalisation
Table 2 displays the article level classification results for the various generalisation splits outlined in Sec. 4. The larger the deviation between the in distribution (ID) articles in the validation set and the out-of-distribution (OoD) articles in the test set, the worse we consider the model’s generalisation performance.
Firstly, we note that classification performance falls short of desired. While the F1-score for the reliable+class tends to be high (in the range of $0.85-0.95$ at a $\approx$60class proportion), classifying unreliable-articles is considerably more difficult—a trend that holds consistently across generalisation forms. This is largely due to low recall scores for the unreliable-class.
Both ‘Publisher’ and ‘Topic’ splitting show moderate decreases in MCC scores, carried primarily by a decrease in the F1-scores for the unreliable-class. Generalisation to completely unseen publishers or topics, cases where one would expect distinctly different linguistic style or vocabulary, is more challenging. The magnitude of this performance degradation, however, is smaller than we initially expected. We attribute this to two effects:
  1. 1.
    More mainstream, prolific publishers are obscuring performance on publishers with fewer articles (see App. C). We correct for this effect by including a publisher-level analysis in Sec. 6.1
  • 2.
    The training data is heterogeneous enough for the models to learn generalisation across publishers. We test for this in Sec. 6.2
  • Since it is conceivable that different publishers prefer particular topics, we compute a correlation degree between the produced test sets. While we find a small but consistent overlap between the two, we do not believe this alone accounts for the similarity in performance (see App. A.2).
    We see no degradation in performance when applying the model to articles from an unseen event (here, the COVID-19 pandemic). Despite the introduction of many unseen terms to the articles’ vocabulary, it appears the manner in which established publishers discuss this new event deviates little from preceding articles.
    The final two generalisation axes exclude a particular misinformation type or political bias from the training set. For the former, we can see little to no effect when removing the ‘Conspiracy-Pseudoscience’ class of articles, but a drastic one if removing ‘Questionable Source’ articles. We posit this is due to the ‘Questionable Source’ being the class of articles written with the explicit purpose of mimicking reliable+publishers, whereas ‘Conspiracy-Pseudoscience’ tends to discuss completely separate topics. In other words, the conspiracy or pseudo-scientific articles tend to be easier to identify as unreliable-.
    For the ‘Political Bias’ generalisation, we see an inability to generalise to opposing political biases. Training on centre and right biased articles sees a $0.19$ drop in MCC, whereas training on centre and left yields a drastic $0.37$ drop. While this is a form of publisher splitting, in both cases the magnitude of the degradation is substantially larger. Especially for transfer to right-biased articles, there exists a drop for both reliable+and unreliable-classification, indicating that it is more challenging for the model to determine article reliability.
    5.2 Generalisation across time
    {tblr}
    colspec=cccccccc, columns = colsep=4pt, cell33 = bg = lightgray, cell44 = bg = lightgray, cell55 = bg = lightgray, cell66 = bg = lightgray, cell77 = bg = lightgray, cell88 = bg = lightgray
    Table 3: Article level MCC scores of models trained with uniform splitting on different years of the dataset. Models are evaluated only on publishers also seen during fine-tuning.
[…]
    When applying the models to unseen years, we find the models to be surprisingly robust, as shown in Table 5.2. At the article level, despite consistent degradation in performance, proximal years tend to achieve similar scores. Only on very distant years does performance degrade dramatically.
    We speculate that these differences are due to differences in the various dataset iterations, while publisher style or idiosyncrasies being relatively static. For example, all models not trained on the 2017 iteration perform poorly on the 2017 iteration (between $0.26$ and $0.34$ MCC), whereas the 2020–2022 editions perform reasonably well on each other’s years. Indeed, visually, Table 5.2 correlates strongly with Appendix C Table B.6, measuring the amount of overlap in publishers across dataset years.
    5.3 LLM Performance
    We compare the performance of the fine-tuned models to that of llama-3-8b-instruct [^6] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct Llama Team (2024), prompted to determine reliability of an article in a 0-shot setting (see App. E.1). We allow the model to ingest the first 512 tokens of the article [^7] Matching the amount of context the fine-tuned models receive.
    In Table 4 we compare the MCC score of the LLM to the 6 ‘Uniform’ fine-tuned models, evaluated on seen and unseen years. Despite the LLMs parameter count and the recency of its pre-training data, we find LLMs to be worse news reliability classifiers than the fine-tuned models. With the advent of LLM-based search engines, this is a somewhat worrying result, and aligns with recent results Liu et al. (2023); Yang and Menczer (2024).
    It should be noted that the use of even this modestly sized LLM incurs a computational cost far greater than fine-tuning. In our experiments, using a single A100 GPU, the LLM took ~$70$ hours to yield a prediction for all articles in the corpus, whereas each of the fine-tuning models were trained and evaluated on all remaining data in ~$12.5$ hours.
    Model MCC
    LLaMA-3 8B Instruct 0.25
    DeBERTa-v3 Unseen Year 0.39
    Seen Year 0.46
    Table 4: Comparing article-level MCC performance of an LLM classifier in a 0-shot setting to fine-tuned models.
    6 Analysis
    6.1 Determinants of Performance
    The analysis of our results, thus far, has been constrained to article-level classification. While this reflects how misinformation classifiers interact with articles, it does not match how we annotate the dataset and can obscure performance on smaller, less mainstream publishers. Ideally, as highlighted by Baly et al. (2018b); Burdisso et al. (2024), classification performance is also evaluated at the publisher-level, testing which publisher properties aid or interfere with misinformation detection. To that end, we employ a binomial logistic regression on the average publisher-level accuracy to assess which aspects of a publisher determine the achieved accuracy score (for details and full model specification, see Appendix E.4). Unlike standard logistic regression, the dependent variable represents the ratio of correct classifications to all articles for each publisher individually (i.e., samples from a binomial distribution).
    In Figure 1 we show coefficient magnitudes for several important determinants, expressed as effect sizes Chinn (2000); Lampinen et al. (2022). A positive effect size indicates that the variable increases the odds of accurate classification, ceteris paribus. While interpreting the effect size magnitudes, the variable’s range of possible values must also be considered (see Appendix E.4 Table 16).
    The size of the training set has a large positive effect, with a $1.91$ multiplicative increase in the odds for each 10-fold increase in training samples. Thus, a publisher with 1000 articles in the training set is $3.82$ times more likely to have correct classification in the test set than a publisher with only 10 articles. Foreign publishers also prove easier to identify.
    Figure 1: Coefficients of the determinants model, expressed as effect sizes. Circles are centred about the effect size, with lines giving the 95% confidence interval.
    Relative to centre-biased publishers, publishers on the left or right sides of the political spectrum see slightly degraded performance, even after controlling for the publisher label. Moreover, classification on the unreliable-classes suffers more than on reliable+publishers. Since the odds ratios or effect sizes are difficult to interpret, we also provide the estimated marginal mean publisher-level accuracy for those combinations in Table 6.1.
    Despite right biased unreliable-sources being far more prevalent in the training data (relative to left and centre bias), for both the ‘Questionable Source’ and ‘Conspiracy-Pseudoscience’ classes, their performance is worse. This somewhat confirms the results in the ‘Political Bias’ rows of Table 2: models struggle disproportionately to discriminate between reliable and unreliable-right-biased articles.
    We see a slight positive correlation with the MBFC ‘Conspiracy-Pseudoscience’ level. The higher the value, the further the publishers’ articles tend to deviate from convention. As a result, strongly conspirational or pseudo-scientific publishers are $4.84$ and $4.72$ times more easily identified than publishers where this effect is weaker.
    Finally, when looking at the factuality level (the propensity for a publisher to publish factual articles) we find a positive interaction for reliable articles, and weakly negative interactions for unreliable articles. The more a publisher goes against the expectation (factual for reliable for publishers, false for unreliable), the more difficult it becomes to disambiguate the source.
    {tblr}
    colspec=rcccc, columns = colsep=2pt,
    Table 5: Predicted publisher-level accuracies averaged over combinations of the MBFC label (columns) and political bias (rows).
[…]
    6.2 Effect of Publisher Diversity
    Finally, we test to what extent the diversity of the publishers present in the training data has an effect on the generalisation capacity of the models. We re-run the ‘Publisher’ split experiment with smaller training sets, constrained to only the most prolific publishers. Specifically, for each MBFC label, we only include the top-n most frequent publishers in the training data, while leaving the test set untouched. While this reduces the amount of variation in publishers considerably, it minimally affects the total amount of data present. Each training set still consists of hundreds of thousands of articles.
    Fig. 2 displays the generalisation gap (in terms of MCC) induced when increasing publisher homogeneity.
    Especially when limiting performance to the Top-1 most common publishers, the models show increased overfit to the training set. Where the ‘Publisher’ split saw a $0.1$ MCC delta, this increased to an average degradation of $0.5$. As the number of included publishers increases, the generalisation gap decreases and starts to converge to the previously seen ‘Publisher’ values. Notably, the variance in values is also substantially higher in the limited publisher settings.
    From this, we conclude that 1) the splitting described in Sec. 4 to have minimally altered the heterogeneity present in the dataset, and 2) the models improve with publisher heterogeneity.
    Figure 2: MCC scores for different ‘Publisher’ split test sets with only the top-1,2 or 3 most prolific publishers retained. The lightly shaded circles provide a value for each year of the dataset, the solid circles their average. The line in between provides the magnitude of the difference in performance across the seen and unseen publishers.
    7 Conclusion
    This paper introduces misinfo-general, a dataset designed for training and evaluating out-of-distribution generalisation in misinformation detection models. We have defined and operationalised various axes of generalisation, some of which align with prior work, and some that are novel. We have shown that this dataset is challenging for models, and especially so when required to generalise to unseen forms of content or publishers.
    Leveraging the metadata annotations available to us, we further analysed determinants of model performance, finding large discrepancies across political bias and misinformation forms, and have tentatively shown that increased diversity has a positive effect on generalisation.
    While publisher-level labelling introduces noise, we believe the increased scale, diversity, and affluence of metadata make up for this. The first two properties can enable OoD generalisation or robustness, whereas the latter enables evaluation, analysis and potentially generalisation-aware training.
    We make the dataset publicly available, and hope it will serve as a resource for OoD generalisation focused training and evaluation. We hypothesise, but leave as an open research direction, that this large dataset might serve as a pre-training step to boost performance on smaller down-stream misinformation corpora. Otherwise, the use of publisher-level metadata might serve as a test-bed for generalisation focused misinformation detection algorithms.
    8 Limitations
    While the dataset includes a diverse set of publishers, events and topics, ultimately the publisher metadata comes from a single source. The information provided by MBFC assumes a narrow, US-centric world-view. This is especially prominent when discussing foreign publishers. As such, the metadata provides limited nuance, providing only a single sample for an inherently subjective assessment. It is expected that across cultural backgrounds, the publisher information is bound to change.
    In general, the news included in the dataset is US-centric, with the vast majority of publishers being American, producing articles for an American audience. This is exacerbated by us excluding all non-English articles. This means cross-lingual or cross-cultural generalisation cannot be evaluated.
    That said, we do discuss one weak form of cross-cultural transfer. Besides prioritising different events, the primary distinction between the political ideologies discussed in Sec. 4, are their differing norms and values. As such, the poor political bias generalisation bodes ill for the more general cross-cultural generalisation tasks.
    Specific to our approach, we experimented with only a single publisher to class label mapping. By binarising the labels, we leave no room for interpretation by the model. It is apparent that not all unreliable sources are equally unreliable, as shown in Sec. 6.1. Similarly, reliable sources can produce sensational content. Prior work has used different mappings, and we suggest future approaches to test various labelling schemes. With the provided metadata, we believe this can be easily implemented.
    9 Ethical Considerations
    9.1 Dataset Access and Licensing
    As mentioned, despite initial releasing the datasets under a CC0 1.0888https://creativecommons.org/publicdomain/zero/1.0/deed.en, the authors have de-accessioned their original repository since January 1st, 2024. Upon request, the authors note their desire to restrict usage to non-commercial research.
    While we want to respect the wishes of our colleagues, we deem the needs of the research community to outweigh these concerns; without access to misinformation texts it is not possible to develop methods for mitigating their effect.
    Despite this, we (plan to) release our dataset under the more restrictive CC BY-NC-SA 4.0999https://creativecommons.org/licenses/by-nc-sa/4.0/deed.en licence. This allows for redistribution and adaption as necessary for academic research, while preventing commercial use-cases and requiring adaptions to maintain these restrictions.
    Furthermore, we (plan to) release the dataset through two media that allow for restricted access. Specifically, we (plan to) use Harvard’s Dataverse (which implements per-file access restrictions) and HuggingFace’s Dataset Hub (which implements repository-level gating). We initially plan to review access applications manually, limiting use-cases to academic research only.
    9.2 Content
    This dataset inherently contains texts that are toxic, hateful, or otherwise harmful to society if disseminated. The dataset itself or any derivative formats of it, like language models, should not be released for non-research purposes.
    The texts themselves might further be copyrighted by their original publishers. We have ensured that all texts have been ’poisoned’ with tokens that occlude large parts of the text, extending the efforts of the original NELA authors.
    We have deliberately removed all social media content, and all hyperlinks to such content. We consider such content Personally identifiable information (PII), with limited use in misinformation classification beyond author profiling. Such applications are fraught with ethical problems Mishra et al. (2021), and likely only induce overfitting in text-based classification.
    Appendix A Additional Information on Generalisation Axes and Splits
    1. 1.
      Uniform: for this generalisation axis, we employ the industry standard method of stratified and shuffled train/validation/test splits. We maintain the ratios 80/10/20, respectively, throughout. This is meant to form the baseline set of values, both in terms of ID-OoD articles performance, but also the expected delta between those.
    2. 2.
      Time: misinformation changes quickly, but models have been found to be brittle to articles outside the time-span of the training data Bozarth and Budak (2020); Horne et al. (2020b). We train models on a dataset from a single year, and evaluate performance on all other years. To avoid conflating ‘Time’ and ‘Publisher’ results, we only evaluate on articles from publishers seen in the training set. The proportion of publishers–and as a result, the reliable/unreliable proportions–does change drastically across dataset years.
    3. 3.
      Event: we define events as occurrences that span news articles with a definite time-span and sudden occurrence. Novel events spawn articles with a markedly different vocabulary, introducing words, names, or terms not yet available to the model. Prior work has shown models suffer when evaluated on unseen events Lee et al. (2021); Ding et al. (2022). The dataset contains a multitude of such events, but we focus specifically on the COVID-19 pandemic; an event that particularly notable for its rapid onset and the volume of misinformation it spawned Bradd (2024). Models are trained on a subset of articles not containing COVID-19 keywords, and evaluated on ones that do. The article discovery-process is outlined in more detail in Appendix A.1.
    4. 4.
      Topic: we define topics to be relatively static collections of events, wherein the style or publishers’ opinions change little across years or decades. We discover such topics automatically, bottom-up. Specifically, we apply a heavily modified variant of the BERTopicGrootendorst (2022) algorithm to discover the many thousands of events that occur in each year of the dataset. Each event is textually represented by a TF-IDF vector. To group events into topics, we apply a second round of spectral clustering on the adjacency matrix induced by the inter-event cosine distance matrix. The number of articles contained by a topic differs substantially (see Table 7). We reserve the $n$ smallest topics, which collectively contain roughly 20% of all articles for testing, and train on articles from all other topics.
    5. 5.
      Publisher: all scraped articles are annotated with the news website, outlet, or publisher that produced the article (see Fig. 4). In general, news publishers have some editorial bias or stance, making their corpus of articles distinctly different from that of other publishers. Prior work has found misinformation classifiers to be particularly sensitive to changes in publisher Barrón-Cedeño et al. (2019); Bozarth and Budak (2020). We specifically test for this by reserving the $n$ smallest publishers, which collectively contain roughly 20% of all articles for the test set, and train on articles from all other publishers. We stratify this splitting across political bias, to ensure each political bias is represented equally in both the train and test sets.
    6. 6.
      Political Bias: separate from editorial bias, publishers tend to exhibit a political bias as well. MBFC has annotated all publishers on a left-right political bias continuum. We map all ‘Extreme Left’ and ‘Left’ publishers to a coarser ‘Left’ group, and vice versa for the ‘Right’-side of the political spectrum. All other political biases (‘Center-Left’, ‘Center-Right’, ‘Least Biased’ and ‘Pro-Science’) were mapped to ‘Center’ bias. We reserve either the ‘Left’ or ‘Right’ group for evaluation, and train on the collection of articles from the opposite political bias along with all ‘Center’ biased publishers.
    7. 7.
      Misinformation Type: misinformation presents in various forms. MBFC divides unreliable publishers in three categories: ‘Questionable-Source’ (publishers that exhibit extreme bias or propaganda), ‘Conspiracy-Pseudoscience’ (publisher aligning with known conspiracies or pseudo-scientific practises) and ‘Satire’ (publishers whose content is purposefully false for comedic effect). To test this form of generalisation, we reserve either ‘Questionable-Source’ or ‘Conspiracy-Pseudoscience’ labels for evaluation, and train on all other articles. The test set includes a stratified sample of reliable sources as well.
    A.1 Identifying COVID-19 Articles
    Figure 3: The top figure shows the proportion of articles published that day which discuss the COVID-19 pandemic. The solid line provides a LOESS smoothed trend line. The bottom figure provides a Google search interest for topics that fall under Coronavirus disease 2019. The different years (2020, 2021, 2022) are displayed as banded vertical columns.
    For ‘Event’ splitting, we focus on a single event present over the latter 3 dataset years: the COVID-19 pandemic. We first derive a set of COVID-19 keywords (e.g., ‘sars-cov-2’, ‘lockdown’, ‘mask’), and combine it with the set of keywords defined by Gruppi et al. (2021). We include all articles from 2020-2022 that include any of these terms in the held-out test set. While this is bound to induce a large number of false positives, it ensures no COVID-19 related terms contaminate the models’ learned vocabulary. In Fig. 3 we display the correlation between the found number of articles and the amount of Google search volume.
    A.2 Test Set Correlations
    Year IoU IoU Over Random Test Set Sizes
    Publisher Topic
    2017 0.1183 9.07% 83633 82688
    2018 0.1227 5.62% 352385 375596
    2019 0.1210 6.80% 459787 483084
    2020 0.1386 24.24% 815825 813176
    2021 0.1317 24.82% 843197 832201
    2022 0.1318 32.93% 821905 807449
    Table 6: The amount of article_id overlap between the ‘Publisher’ and ‘Topic’ test sets. Column ‘IoU’ gives the intersection over the union between the two held-out test sets. Column ‘Over Random’ compares this to the achieved value for 10000 simulated draws from a uniform distribution of the same sizes as the actual test set sizes (given in the last two columns).
    Publishers have a natural tendency to prioritise different topics. This is especially prevalent for less mainstream publishers, which have far fewer resources to spend on the breadth of their reporting. It is therefore plausible that the ‘Publisher’ and ‘Topic’ splits have some overlap.
    We test for this by comparing the amount of article level overlap in the held-out test sets of the ‘Publisher’ and ‘Topic’ to random draws of the same size from uncorrelated uniform distributions. The larger this overlap, the more correlated the two test sets are. We display the results of our test in Tab. 6. While we find more overlap than random draws, the effect is not large. Even in the most egregious case, there is only a $0.04$ difference in absolute IoU. As such, we believe the two different test sets measure distinct generalisation aspects.
    Statistic 2017 2018 2019 2020 2021 2022
    Article Counts 0.10M 0.46M 0.59M 1.02M 1.04M 0.99M
    Label Proportion Reliable 61.39% 72.82% 75.91% 70.73% 71.36% 72.00%
    Unreliable 38.61% 27.18% 24.09% 29.27% 28.64% 28.00%
    Publishers Reliable 43 99 134 189 183 184
    Unreliable 49 60 100 249 219 201
    Truncated Token Counts Mean 398.53 383.59 430.22 432.87 428.15 435.90
    Quantiles 299, 488, 512 266, 448, 512 358, 512, 512 369, 512, 512 366, 512, 512 383, 512, 512
    Event Clustering # Events 2674 6288 7718 7931 8758 8336
    Mean Event Size 38.67 73.38 76.87 128.49 118.24 119.29
    Smallest Topic Size 2.7K 20.5K 24.6K 27.7K 51.3K 48.4K
    Largest Topic Size 38.3K 94.1K 113.1K 192.3K 211.4K 209.4K
    Reliable Labels Left 17.91% 13.32% 11.13% 9.01% 8.76% 9.88%
    Left-Center 24.09% 29.62% 35.19% 36.50% 34.53% 35.01%
    Least Biased 0.93% 2.18% 4.38% 4.91% 5.13% 3.97%
    Right-Center 4.34% 12.66% 9.81% 8.70% 10.38% 9.52%
    Right 14.12% 15.05% 15.41% 11.40% 12.03% 13.08%
    Pro-Science 0.21% 0.53% 0.55%
    Unreliable Labels Satire 1.70% 0.99% 0.82% 0.66% 0.48% 0.42%
    Questionable Source 27.83% 20.14% 17.34% 19.41% 19.00% 18.45%
    Consp.-Pseudosci. 9.09% 6.05% 5.93% 9.20% 9.17% 9.13%
    Table 7: Various statistics on the dataset and labels. Article Counts provides the total number of articles in each dataset year. Label Proportion provides the relative occurrence of reliable vs. unreliable articles, whereas Publishers provides the number of such publishers present. Section Truncated Token Counts provides the mean and the 25, 50, 75th quantiles of the number of tokens per article. Note that is after truncation at 512; the raw articles tend to be much longer. Event Clustering provides the number and average size (in articles) of events. We also provide the number of events belonging to the smallest and largest topics. Sections Reliable Labels and Unreliable Labels provides the proportion of each MBFC label in the dataset, split into their class label.
    Appendix B Dataset Processing & Statistics
    Year 2017 2018 2019 2020 2021 2022 Total
    Original 0.14 0.70 1.12 1.78 1.86 1.78 7.24
    De-aggregation & Labelling 0.13 0.61 0.96 1.62 1.71 1.66 6.69
    -1% -9% -12% -5% -3% -3% -8%
    Exact Deduplication 0.12 0.55 0.86 1.34 1.39 1.32 5.58
    -6% -11% -12% -18% -20% -21% -23%
    Cleaning 0.12 0.53 0.71 1.12 1.16 1.11 4.74
    -3% -4% -17% -16% -17% -16% -35%
    Near Deduplication 0.11 0.47 0.65 1.03 1.06 1.01 4.33
    -11% -12% -8% -8% -9% -9% -40%
    Language Detection 0.11 0.47 0.64 1.02 1.05 0.99 4.16
    0% -1% -1% -1% -1% -1% -43%
    Table 8: Size of the datasets, in terms of millions of articles, after each step of cleaning, per year. The lower percentage gives the step-to-step reduction in size. The ’Total’ column computes reduction relative to ’Original’.
    We downloaded the initial NELA corpora from the Harvard Dataverse [^10] The corpora have since been de-accessioned, and can longer be downloaded. We expand on this in Sec. 9, under a CC0 1.0 licence [^11] https://creativecommons.org/publicdomain/zero/1.0/deed.en.
    Publisher # Articles
    drudgereport 92 186
    bonginoreport 17 895
    whatreallyhappened 57 431
    thelibertydaily 6 346
    yahoonews 52 797
    theduran 10 722
    Total 237 377
    Table 9: Publishers removed for being news aggregation sites, with article counts post de-aggregation.
    Domain # Articles
    soundcloud.com 106
    youtu.be 4 009
    apps.apple.com 414
    amazon.com 112
    facebook 108
    amazon 113
    youtube 2 902
    play.google.com 421
    twitter 274
    instagram 15
    reddit 4
    dnyuz.com 1 900
    Total 10 378
    Table 10: Publishers removed for being social media sites, with article counts post de-aggregation.
    B.1 Re-Labelling
    As a first processing step, we relabelled all publishers. This was done to 1) attribute articles to their original publisher (where possible), 2) ensure publisher information was up-to-date (MBFC had expanded their catalogue considerably) and 3) to mitigate the effects of publishers that might interfere with the learning signal.
    An important class of publishers that belong under that third point are aggregation sites. Such sites either do not produce original content, or intersperse articles from (usually more reputable) other sources through their content. While the collection of articles as a whole might express some editorial bias, for the most part, these sorts of publisher introduce noise into an already noisy labelling scheme.
    We manually re-mapped all URL domains to a set of publishers consistent across years, excluding all news aggregation platforms and social media sites (see Tables B and B).
    It should be noted that the 2018 edition of NELA did not contain URLs, making relabelling considerably more difficult.
    Table B.6 shows the amount of publisher overlap exists between different dataset years.
    Type Sample Text
    Banners Click for more article by Guest …
    Previews To read the full blog, please check out the complete post on <source>
    Prayers Our Father, who art in heaven, hallowed be thy Name …
    Podcast Descriptions Don’t forget to tune in to …
    Error Messages 403 Forbidden nginx
    Table 11: Frequent duplicated articles.
    Rule VALUE
    Articles must contain at least \${VALUE} tokens 16
    Articles must not contain more than \${VALUE} tokens 4096
    The article must have a title -
    The article must be at least \${VALUE} times longer than its title 3
    The article must have a mean token length greater than \${VALUE} characters 2
    The article must have a mean token length less than \${VALUE} characters 10
    The article must have at most \${VALUE}copyright tokens 20
    Table 12: Filtering rules.
    B.2 Deduplication
    Models trained on this dataset have to (implicitly) learn a mapping from an article to its publisher’s class. As a result, any instances of duplicate content unique to a publisher or publisher class likely induces overfit. The model need only memorize those cases to make a prediction, independent of the article content.
    To minimize the effect of these duplicates, we apply several stages of deduplication.
    1. 1.
      we only keep the first (as determined by publication date) instance of exact duplicates for each publisher. Besides plagiarism, the most common duplicates are due to errors in scraping or parsing, resulting in artefacts (see Table 11)
  • 2.
    we remove all articles with duplicate titles or URLs, across the entire dataset. This primarily removes articles that are updated at a later stage (e.g., live-blogs or summaries), or result from URL re-directions 3. we remove all articles if more than 5% of its sentences are duplicates from the same publisher
    The former two (‘Exact Deduplication’ in Tab. 8) can handle document-level duplicates, like plagiarised articles or updated blog-posts, but miss near duplicates. The latter should filter out near duplicates, like lightly edited documents, (‘Near Deduplication’ in Tab. 8).
    An added benefit of these deduplication steps is the identification of common scraping errors. In fact, we found this to be the most frequent form of duplication. We list some common texts in Table 11. Since the majority of such errors were automated responses unique to individual publishers, we could filter these out with relative ease using the described ‘Exact Deduplication’ approaches.
    Ultimately, the separate deduplication steps together resulted in removing 3.1M articles, comprising roughly $\approx 22\%$ the dataset after relabelling.
  • B.3 Cleaning & Filtering
    Many of the included articles either include too many non-semantically relevant tokens or are otherwise malformed. This is aggravated by the NELA authors ’poisoning’ the dataset (from 2019 onwards) with repeated ‘@’ tokens to avoid copyright infringement. We do our best to clean the article texts, and remove those with no discernible semantic information.
    We normalize all punctuation and remove any embedded URLs, HTML or Markdown markup. SpaCy’s Montani et al. (2023) en_core_web_sm was used to annotate all tokens in the corpora. We identify self-references as named entities with a large longest common substring relative to the publisher’s name. We also mask any sentences which SpaCy flags as being part of an email, URL or twitter handle. Finally, we standardize the NELA copyright poisoning, applying it to all dataset years equally.
    All-in-all, this introduces 4 new special tokens: <copyright> replacing NELA’s repeated @ tokens, <twitter> for X (formerly known as Twitter) handles, <url> for any URLs and <selfref> for any self-references.
    To ensure all articles contained enough grammatical text to reasonably classify, we removed any article which did not abide by the rules delineated in Table 12. As a final step, we use Lingua Stahl (2024) to filter out any non-English texts.
    Despite removing almost half of the articles, the dataset retains a level of ’noise’ customary to data sourced from the internet. This is inherent to the domain, and further cleaning might negatively affect the realism of the benchmark. One type of noise that could interfere with learning is the presence of stylistically unique substrings identifying publishers. For example, articles from the same publisher tend to contain similar by-lines, attribution messages, or donation requests.
    B.4 Event & Topic Clustering
    To generate the metadata necessary for the ‘Topic’ generalisation form (see Sec. 4 or App. A), we opted for a bottom-up approach. This involved first clustering the dataset into thousands of fine-grained events, before clustering the event clusters into overarching topics.
    To achieve this, we used a heavily modified variant of BERTopic Grootendorst (2022) with a gte-large [^12] https://huggingface.co/thenlper/gte-large Li et al. (2023) backbone.
    Figure 4: Schema for the generated metadata database. ’PK’ indicates primary key, ’FK’ denotes a foreign key relationship.
    After embedding an entire year of the dataset, we first reduce their dimensionality using mini-batched PCA, and whiten the data. We construct a vocabulary over all the documents, by aggregating mini-batches of vocabularies. We then apply UMAP dimensionality reduction, using the PCA solution as initialisation, and the cosine distance as the distance metric. Clustering was performed via HDBScan to mini-batches of the embeddings, assigning all articles to their nearest event cluster. Using the dataset-wide vocabulary, we generate a single TF-IDF representation matrix.
    After mini-batching, we generate a hierarchical clustering on top of the event-based TF-IDF matrix. We merge any events with a distance below 0.8 to mitigate the effect of mini-batching, and re-construct the TF-IDF representations. We deem these the final event representations. Table 7 provides the total number of events and the average number of articles contained within each.
    Finally, we apply a round of spectral clustering to these representations, compressing these events into 10 groups, which we deem topics. The numbers of events in each topic varies (see Table 7).
    B.5 Metadata
    Article content is stored in arrow files, accessed using HuggingFace’s datasets library HuggingFace (2021). We store all generated metadata in a duckdb database Mühleisen and Raasveldt (2024). Fig. 4 depicts the metadata schema. Table ’Sources’ provides information sourced from MBFC on different publishers, whereas tables ’Articles’ and ’Topics’ provides information on the produced articles and events/topics, respectively. The articles are all given unique article identifiers. This ensures we can quickly generate relevant splits of the dataset and link these to article-level predictions.
    B.6 Publisher-level Metadata
    All publisher metadata is stored in the ’Sources’ table. Each source is identified by a unique name, linked in a 1-to-many relationship to articles in the ’Articles’ table. The url column provides a link to the MBFC page with the source metadata. The label column provides the MBFC label, also shown in the lower rows of Table 7. In column bias, the political bias (one of ’Extreme Left’, ’Left’, ’Left-Center’, ’Least Biased’, ’Pro-Science’, ’Right-Center’, ’Right’ or ’Extreme Right’) of the publisher is provided. The factuality column provides an ordinal value for the propensity of a publisher to report factual news. Columns country and type provide information about the country of origin of the publisher, and their primary form of publication (e.g., TV, blog posts), respectively. In column popularity, the average number of visits to the publisher’s main website is provided as an ordinal categorical value. credibility provides an overall assessment of the publishers’ credibility, aggregating all other variables into a single score. Finally, the conspiracy and pseudosci columns provide the ‘strength’ of the conspiracy or pseudo-science source. The larger this value is, the more these sources deviate from the public or scientific consensus. This is only provided for publishers labelled as ‘Conspiracy-Pseudoscience’.
    Figure 5: The authorship of articles in the dataset, aggregated into deciles. The right most column of each panel represents the top 10% most prolific publishers, the left most the bottom 90%.
    {tblr}
    colspec=ccccccc, columns = colsep=4pt, cell33 = bg = lightgray, cell44 = bg = lightgray, cell55 = bg = lightgray, cell66 = bg = lightgray, cell77 = bg = lightgray, cell88 = bg = lightgray
    Table 13: Overlap of publishers between years.
[…]
    Appendix C Publisher Statistics
    Across all dataset iterations, the vast majority of articles are written by a small minority of prolific publishers. This is shown in Figure 5. The most prolific publishers have authored between $40$-$60$% of articles within a year. The next deciles manage only $20$%, then $10$-$15$%, etc., with the smallest publishers authoring only a few articles. This effect seems to get more drastic for the later dataset years.
    As discussed in Sec. 5 this can obscure poor performance when using article-level evaluation metrics. In such cases, the average performance score will tend to be dominated by the performance of the largest publishers, instead of the body of publishers as a whole. As shown in Figure 1 and Tables 16 and 17, performance on these less mainstream publishers tends to be substantially worse. As a result, the OoD performance metrics in Table 2 likely underestimate the generalisation gap.
    This dataset property means publisher-level analyses are necessary, in addition to the far more common article-level performance measures. That said, prominent misinformation datasets tend to sample from a far more a smaller and homogenous set of publishers, with typically more balanced authorship proportions.
    Another noteworthy statistic is the amount of overlap between publishers across the dataset years. While the dataset as a whole contains $488$ distinct publishers, as can be seen in Table 7, each dataset year contains fewer total publishers. Furthermore, the distribution of those publishers (in terms of article counts) shifts across the years. This is partially due to a changing collection methodology used by Horne et al.; Norregaard et al.; Gruppi et al.; Gruppi et al.; Gruppi et al..
    As a result, the different dataset years have different degrees of overlap between each other. Table B.6 shows the publisher intersection-over-union overlaps between the $6$ different dataset years. The largest block of overlap is between the years 2020-2022, with overlaps across datasets being well above $0.80$. The IoU scores to other dataset years are far lower, with 2017 having particularly low overlaps.
    Appendix D Training Details
    Hyperparameter Value
    seed 942
    max_length 512
    val_prop 0.1
    batchsize 64
    pooler_dropout 0.1
    lr
       embeddings 5.0e-7
       pooler 5.0e-5
       classifier 1.0e-3
    weight_decay 0.1
    lr_scheduler
       type ‘polynomial’
       power 3.163
    max_steps 3.0e+6
    warmup_ratio 0.0213
    eval_prop 0.05
    patience 2
    Table 14: An overview of important hyperparameters and their values. These were tuned on the validation set of the ‘Uniform’ split dataset.
    While DeBERTa-v3 models can theoretically handle infinite sequences, to constrain memory requirements we truncate tokenization after the first $512$ tokens. This allowed for a batch size of $64$ during training, and $512$ during validation. We used Adam with weight decay Kingma and Ba (2017); Loshchilov and Hutter (2019) as the optimizer.
    Some other important training hyperparameters are listed in Table 14. We set a high learning rate for the classifier and pooler, but a very low value for the embeddings layer to avoid catastrophic forgetting. We employed a polynomial decaying learning rate scheduler, which first linearly warms up, before decaying according to η_t=(tmax_steps)^powerη_t-1.. For each experiment, regardless of the size of the dataset, we allow for a maximum of $63.0e+$ 6 update steps, evaluating on the validation set every $5$%. We employ early stopping with patience of 2, through which training usually concludes well before reaching the maximum budget.
    D.1 Additional Training Details
    All experiments were conducted on a Linux SLURM-based academic cluster. Nodes consisted of an Intel Xeon Platinum 8360Y CPU with 18 cores, 2.4 GHz, a single NVIDIA A100 GPU accelerator (40 GiB of HMB2 memory) and 128 GiB of DDR4 memory.
    We use Python 3.11 and PyTorch 2.2.2 built with CUDA 11.8. For training, we make extensive use of utilities implemented in HuggingFace’s transformers library: transformers 4.37.2, datasets 2.19.0 and accelerate 0.30.1. All experiments were conducted under random seed 942. For local development, we use Ubuntu 20.04.6 LTS (GNU/Linux 5.15.90.1-microsoft-standard-WSL2 x86_64).
    Figure 6: Probabilities of making an error for publishers of a particular label and political bias, relative to the probability of making a mistake for publishers of the same label. FP columns denote the probability of error for reliable publishers, and FN vice versa. The higher
    Appendix E Additional Results
    E.1 LLM Inference
    To conduct LLM inference, we use meta-llama/Meta-Llama-3-8B-Instruct using the following prompt:
    <s>[INST] <<SYS>> You are a content moderator working with journalistic articles. Your task is to identify articles from unreliable publishers. <</SYS>>
    Does the following text come from a reliable news publisher? Respond with ’yes’ or ’no’.
    Article: \${ARTICLE}
    Does this article come from a reliable news publisher? ’yes’ or ’no’: [/INST]
    Here, \${ARTICLE} is replaced by the first $512$ tokens of each article.
    We use the pmi_dc decision rule, introduced by Holtzman et al. (2021). We use the same prompt without the article tokens as the domain conditional text. Empirically, this performs slightly better across the entire corpus.
    Due to the (very) long articles prevalent in this dataset, using few-shot exemplars was deemed intractable.
    E.2 Political Bias
    To further assess the model’s interaction with political bias, we compute the complement of the expected publisher-level accuracy score for a political bias and publisher label (reliable or unreliable), and divide it by the marginal expected score:
    $\frac{1-\mathbb{E}[\text{{Acc}}_{p}|\text{bias}(p),y_{p}]}{1-\mathbb{E}[\text{{Acc}}_{p}|y_{p}]}-1,$
    essentially giving the shifted exponentiated pointwise mutual information (PMI) of making an error for a particular combination of publisher political bias and label,
    $\exp\quad$ $\text{pmi}\left(\hat{y}_{p}\not=y;\text{bias}(p)|y\right)-$
    =$\frac{p(\hat{y}_{p}\not=y|\text{bias}(p),y)}{p(\hat{y}_{p}\not=y|y)}-1.$
    The higher this value, the greater the association between making an error for a particular political bias deviates from the expected value. If an error is made for a reliable publisher, this corresponds to a False Positive (FP) whereas an error for an unreliable publisher would correspond to a False Negative (FN). We introduce the shifting to ensure no bias corresponds with 0, whereas positive and negative bias correspond to positive and negative values, respectively.
    In Figure 6 we display various such biases for each iteration of the corpus. As seen in Table 16, performance degrades for the more extreme political biases. There is, however, a notable difference between left-biased and right-biased publishers, and especially for the reliable right-biased publishers. The probability of the model generating a False Positive is substantially higher (sometimes near twice as likely) than for reliable leftist or centre biased publishers. We conclude that the model tends to confuse articles from reliable right biased publishers with unreliable ones, far beyond what would be considered ’due to chance’.
    On the other hand, the model tends to be overly optimistic for centre-biased publishers, with lower FP probabilities and a tendency for slightly higher FN probabilities. An unreliable publisher could therefore escape moderation by limiting its explicit political views. While this is a relatively unlikely example, as most unreliable publishers are unreliable because of overt extreme political views, it does showcase an important blind-spot for these models.
    Given the high-stakes nature of misinformation moderation decisions, this indicates the presence of undesirable behaviour. In their current state, our misinformation detection models discriminate against publishers of a particular political group.
    E.3 Publisher Heterogeneity
    In Sec. 6.2 we tested for the effect of publisher heterogeneity on generalising to unseen publishers. We operationalized this by re-running the ‘Publisher’ split experiment with only the top-n most prolific publishers for each MBFC label, while leaving the test set untouched.
    This dramatically reduces the amount of variation in publishers, while minimally affecting the total amount of data present. This can be seen in Table 15. Note that these are expressed as percentages, but in absolute terms constitute hundreds of thousands of articles.
    Quantity Baseline Top 1 Top 2 Top 3
    Articles 100% 28% 43% 55%
    Publisher 100% 11% 20% 30%
    Table 15: Mean number of articles and publishers retained in the training set after limiting to the top-n most prolific publishers. Note that the number of articles and publishers deviates from the distributions depicted in Fig. 5. In this case, we additionally condition on the publisher’s MBFC label, across which publishers are not equally productive (see Table 7).
    E.4 Determinants of Publisher-Level Accuracy
    To estimate which factors impact accuracy at the publisher level. We define publisher level accuracy for a publisher $p$ as,
    $\text{{Acc}}_{p}=\frac{1}{|\mathcal{D}^{(p)}|}\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)}),$
    or put otherwise, the expected rate of correct classification for all documents belonging to that publisher.
    To determine which factors contribute to these scores, we use a multinomial logistic regression model. Specifically, we use statsmodels Seabold and Perktold (2010) to estimate the dependent variables as,
    $\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)})+\mathds{1}(\hat{y}^{(d)}\not=y^{(p)})$
    $=\sigma(\beta_{0}+\sum_{i=1}^{K}\beta_{i}x_{i}),$
    We use $\mu^{2}\cdot(1-\mu)^{2}$ as the variance function, Pearson’s $\chi^{2}$ as the scale value and the logit as the link function $\sigma$.
    E.4.1 Uniform Generalisation
    The fully specified model is displayed in table 16, with some important variables’ coefficients depicted graphically in Fig. 1. The left most column provides natural groupings of the different variables. The interpretation of the coefficient magnitudes should be adjusted to consider the range of possible values for the corresponding variable.
    In the Generic category, we include the logarithm of the number of articles present in the training data, and whether the publisher is foreign or not. Both variables are assigned large coefficients. Every 10x increase of labelled articles increases the odds of correct classification by a factor of $1.91$. Foreign publishers tend to be $2.91$ times as easy to classify relative to U.S. based publishers.
    The Year category of variables includes a dummy variable for each iteration of the dataset. Here we find relatively little difference, despite statistical significance, with coefficients falling between $0.85$ and $1.15$.
    We further include dummy variables for the different political biases and MBFC labels (assumes Centre-Reliable to be the default group). We furthermore include an interaction effect between all political biases and MBFC labels. We find that unreliable publishers are substantially more difficult to classify, with all unreliable labels having large negative coefficients. This is especially true for the ’Questionable Source’ category of articles, with log-odds shifts of $-1.25$ and $-1.84$ for left and right biased publishers, respectively. Furthermore, corroborating the analysis in Appendix E.2, the models perform much more poorly for publishers on the right side of the political spectrum.
    In the Strength block of variables, we include MBFC’s strength of conspiracy or pseudo-science. All ‘Conspiracy-Psuedoscience’ sources are rated on a scale, with larger values indicating a further deviation from the societal norm. Reliable sources a given a default score of 00. We find that there exists a slight, but consistent, positive correlation with the strength of the pseudo-scientific or conspiratorial claims. In other words, the more those publishers deviate from the status quo, the more easily these are identified.
    Similarly, for the Factuality Score block, we include the MBFC score for a publisher’s propensity for factual reporting as interaction effects with each MBFC label (excluding ‘Satire’). Again, we find a slight positive correlation for reliable sources, meaning that the more often such publishers report factual information, the better classification performance. For the unreliable publishers, however, the opposite holds: the less likely to present factual news, the easier the classification. In either case, an ’average’ level of factuality tends to correspond to more ambiguous cases, lying closing to the model’s decision boundary.
    E.4.2 Topic Distance
    We repeat this analysis, but this time aggregating entities at the event/label level, using the ‘Topic’ split results. The model specification is left as above, with the introduction of a single additional variable: the minimum cosine distance to any event in the training set. The coefficients are provided in Table 17.
    While comparison across Tables 16 and 17 is not directly possible, as each estimates using different entities and different models, it is encouraging to see similar magnitudes for most variables’ coefficients.
    Surprisingly, we observe a very small but positive effect: $\beta_{\text{min\_dist}}=0.15$. This implies that a topic that is as far away as possible from all topics in the training set sees a $1.16$ increased odds of correct classification. While this is a weak effect, it does suggest that events more typical to a certain topic are easier to classify in than those near the topic cluster borders.
    Group Parameter Range $\beta$ $\exp(\beta)$ Std. Err. p 95% CI
    Intercept Intercept 1 -0.85 0.43 0.03 0.00 -0.91 -0.79
    Generic $\log_{10}(\text{train count})$ $[0,\infty]$ 0.65 1.91 0.01 0.00 0.64 0.66
    Foreign 0..1 1.07 2.91 0.01 0.00 1.06 1.08
    Year 2018 0..1 0.07 1.08 0.02 0.00 0.04 0.11
    2019 0..1 0.14 1.15 0.02 0.00 0.10 0.17
    2020 0..1 -0.16 0.85 0.02 0.00 -0.19 -0.13
    2021 0..1 -0.06 0.94 0.02 0.00 -0.09 -0.03
    2022 0..1 0.00 1.00 0.02 0.85 -0.03 0.04
    Political Bias Left 0..1 -0.39 0.68 0.01 0.00 -0.41 -0.37
    Right 0..1 -0.64 0.53 0.01 0.00 -0.66 -0.62
    Label Conspiracy-PseudoScience 0..1 -0.32 0.73 0.22 0.15 -0.76 0.11
    Questionable Source 0..1 -2.40 0.09 0.02 0.00 -2.44 -2.37
    Satire 0..1 -1.58 0.21 0.18 0.00 -1.94 -1.23
    Interactions Left : Conspiracy-PseudoScience 0..1 -1.65 0.19 0.19 0.00 -2.02 -1.28
    Left : Questionable Source 0..1 1.54 4.69 0.03 0.00 1.48 1.61
    Left : Satire 0..1 -2.15 0.12 0.20 0.00 -2.54 -1.75
    Right : Conspiracy-PseudoScience 0..1 -1.70 0.18 0.19 0.00 -2.06 -1.33
    Right : Questionable Source 0..1 1.20 3.31 0.02 0.00 1.17 1.23
    Right : Satire 0..1 0.56 1.75 0.35 0.11 -0.12 1.25
    Strength Conspiracy 1..5 0.19 1.21 0.02 0.00 0.15 0.22
    PseudoScience 1..5 0.17 1.18 0.01 0.00 0.14 0.19
    Factuality Score Factuality : Reliable 1..5 0.35 1.42 0.00 0.00 0.34 0.36
    Factuality : Conspiracy-PseudoScience 1..5 -0.09 0.92 0.03 0.00 -0.14 -0.04
    Factuality : Questionable Source 1..5 -0.13 0.88 0.01 0.00 -0.14 -0.11
    Table 16: Coefficients of the publisher-level determinants model, using the uniformly split models. Range indicates the possible values each variable can take, with ..absent... . indicating all natural numbers between the extremes (inclusive). $\beta$ provides the logistic parameter, and $\exp(\beta)$ its exponent (i.e. the log-odds ratio). Column ’Std. Err.’ provides the standard error, ’p’ the corresponding $p$-value and ’95% CI’ the confidence interval. In the parameter names, the ’:’ indicates an interaction term.
    Group Parameter Range $\beta$ $\exp(\beta)$ Std. Err. p 95% CI
    Intercept Intercept 1 -0.49 0.61 0.05 0.00 -0.60 -0.38
    Distance Topic Distance [0, 1] 0.15 1.16 0.03 0.00 0.09 0.21
    Generic $\log_{10}(\text{train count})$ [0, $\infty$] 0.56 1.75 0.01 0.00 0.54 0.58
    Foreign 0..1 0.88 2.42 0.01 0.00 0.86 0.90
    Year 2018 0..1 0.23 1.26 0.03 0.00 0.17 0.29
    2019 0..1 0.30 1.35 0.03 0.00 0.24 0.36
    2020 0..1 -0.18 0.83 0.03 0.00 -0.24 -0.13
    2021 0..1 -0.47 0.62 0.03 0.00 -0.53 -0.41
    2022 0..1 -0.27 0.76 0.03 0.00 -0.33 -0.21
    Political Bias Left 0..1 -2.81 0.06 0.47 0.00 -3.74 -1.88
    Right 0..1 -2.58 0.08 0.03 0.00 -2.64 -2.52
    Label Conspiracy-PseudoScience 0..1 -1.51 0.22 0.25 0.00 -1.99 -1.03
    Questionable Source 0..1 -0.54 0.58 0.01 0.00 -0.57 -0.51
    Satire 0..1 -0.73 0.48 0.02 0.00 -0.76 -0.70
    Interactions Left : Conspiracy-PseudoScience 0..1 -0.38 0.69 0.41 0.36 -1.18 0.43
    Left : Questionable Source 0..1 2.11 8.27 0.05 0.00 2.02 2.21
    Left : Satire 0..1 -1.83 0.16 0.31 0.00 -2.44 -1.22
    Right : Conspiracy-PseudoScience 0..1 -0.73 0.48 0.41 0.07 -1.53 0.06
    Right : Questionable Source 0..1 1.92 6.85 0.03 0.00 1.87 1.98
    Right : Satire 0..1 0.32 1.37 0.45 0.48 -0.56 1.19
    Strength Conspiracy 1..5 0.42 1.52 0.04 0.00 0.35 0.49
    PseudoScience 1..5 0.29 1.33 0.03 0.00 0.24 0.34
    Factuality Score Factuality : Reliable 1..5 0.38 1.47 0.01 0.00 0.37 0.40
    Factuality : Conspiracy-PseudoScience 1..5 0.42 1.53 0.05 0.00 0.33 0.52
    Factuality : Questionable Source 1..5 -0.44 0.65 0.01 0.00 -0.46 -0.41
    Table 17: Idem, but now for the publisher-topic aggregated binomial logistic model. Since the aggregation level differs from that of the model presented in Table 16, the coefficients are not directly comparable.